# Amortized Safe Active Learning for Real-Time Decision-Making: Pretrained Neural Policies from Simulated Nonparametric Functions

### v1

@@ C.1.2 Objectives: Unconstrained AL - Regularized Entropy
Here, we discuss empirical issues of unconstrained entropy and our approach using mutual information objectives (Guestrin et al., 2005; Krause et al., 2008).
Maximizing an entropy objective favors a set of distinctive points, which naturally encourages points at the border, as they are the most scattered. Guestrin et al. (2005) argued that exploring the border may be suboptimal because we would rather explore inside the space. The authors further proposed a mutual information acquisition criterion to tackle this problem, at least in conventional AL settings. In our training scheme, the objectives $\mathcal{H}(\phi)$ and $\mathcal{H}_{\text{mean}}(\phi)$ sometimes overemphasize the border and completely ignore the inner region of $\mathcal{X}$, which is clearly not desired. Part of the reasons can be the tanh layer in our NN (Figure A.B.4), which, although helping us bound the NN outputs in $\mathcal{X}$, has vanishing gradients at the boundary. To tackle this issue, we wish to derive a mutual information objective suitable for our framework.
In active GP learning, Guestrin et al. (2005) define a mutual information criterion as the reduction of entropy in an unexplored region: $\mathbb{H}(y(\bm{x}_{\phi,1}),...,y(\bm{x}_{\phi,T})|Y_{\text{init}})-\mathbb{H}(y(\bm{x}_{\phi,1}),...,y(\bm{x}_{\phi,T})|Y_{\text{init}},y(\mathcal{X}\setminus\bm{X}_{\phi}))$, $y(\mathcal{X}\setminus\bm{X}_{\phi})$ means the outputs corresponding to $\mathcal{X}\setminus\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\}$. The original paper considered small discrete $\mathcal{X}$ where $y(\mathcal{X}\setminus\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\})$ is a computable set of random variables. In our framework, $\mathcal{X}$ is a continuous space, and thus this is not well-defined. Even if $\mathcal{X}$ is discrete, conditioning on a large pool (fine discretization) is computationally heavy, i.e. GP cubic complexity $\mathcal{O}\left(|\mathcal{X}|^{3}\right)$ (Appendix A). A discrete pool also enforces a classifier-like policy $\phi$, as the policy needs to select points from a pool, which prohibits us from utilizing the current NN structure developed on top of Foster et al. (2021); Ivanova et al. (2021).
To derive mutual information objectives suitable for our learning framework, note that entropy reduction is a regularized entropy objective. We propose a simple yet effective approach: compute the regularization term only on a sparse set of $N_{\text{grid}}$ samples $\left(\bm{X}_{\text{grid}},Y_{\text{grid}}\right)\subseteq\mathcal{X}\times\mathcal{Y}$. Then we turn the acquisition functions proposed by Guestrin et al. (2005); Krause & Guestrin (2007) into the following training objectives
@@
$\begin{split}\mathcal{I}_{\text{mean}}(\phi)=&\mathbb{E}_{\theta,\sigma^{2}}\mathbb{E}_{f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}}\left[\frac{\mathbb{H}(y(\bm{x}_{\phi,1:T_{\text{sim}}})|Y_{\text{init}})-\mathbb{H}(y(\bm{x}_{\phi,1:T_{\text{sim}}})|Y_{\text{init}},Y_{\text{grid}})}{N_{\text{init}}+T_{\text{sim}}}\right].\end{split}$ (A.15)
$N_{\text{grid}}$ should be much larger than $T$. Maximizing these objectives encourages $\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\}$ to track subsets of $\bm{X}_{\text{grid}}$. The intuition is two-fold: (i) we can view them as entropy objectives regularized by an additional search space indicator, or (ii) we can view them as imitation objectives because a subset of grid points, if happens to have small joint likelihood or large joint entropy, maximizes the objective.
@@ C.2 Objectives: DAD Baseline
0:  2.1, $T$, $N_{\text{init}}$
1:  sample a batch of $(\theta,\sigma^{2})$
2:  for each $(\theta,\sigma^{2})$ do
3:     sample $f_{0,\text{raw}}\sim\mathcal{GP}(0,k_{\theta})$
4:     $f_{0}(\cdot)=f_{0,\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{0,\text{raw}}(\bm{x})]$
5:     sample $\bm{X}_{\text{init}}\sim\text{Uniform}[\mathcal{X}],|\bm{X}_{\text{init}}|=N_{\text{init}}$
6:     sample $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I),Y_{\text{init}}=f_{0}(\bm{X}_{\text{init}})+\mathcal{E}_{\text{init}}$
7:     $\mathcal{D}_{0}=\{\bm{X}_{\text{init}},Y_{\text{init}}\}$
8:     for $t=1,...,T$ do
9:        $\bm{x}_{\phi,t}=\phi(\mathcal{D}_{t-1})$
10:        sample $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),y_{\phi,t}=f_{0}(\bm{x}_{\phi,t})+\epsilon_{t}$
11:        $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t}\}$
12:     end for
13:     sample $f_{l,\text{raw}}\sim\mathcal{GP}(0,k_{\theta}),l=1,...,N_{f,q}$
14:     $f_{l}(\cdot)=f_{l,\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{l,\text{raw}}(\bm{x})],l=1,...,N_{f,q}$
15:  end for
16:  compute DAD loss (Eq. A.16), update $\phi$
We describe the DAD baseline in our GP context (Foster et al., 2021). DAD is a scheme which samples tasks and learns to perform Bayesian optimal experimental design originally aiming at parametric models. Bayesian optimal experimental design aims to query data to learn about a posterior Bayesian model (Rainforth et al., 2024). In a GP learning problem, this means collecting $\bm{x}_{1:T},y_{1:T}$ to make inference $p(f(\cdot)|\bm{X}_{\text{init}},Y_{\text{init}},\bm{x}_{1:T},y_{1:T})$. DAD follows similar training procedure: we (i) sample tasks from a prior, and (ii) learn a data querying policy by optimizing the DAD objective. To train with DAD, one may use our main training algorithm because we generate more quantities than needed by the DAD objective. For clarity, however, we write down a clean version of DAD training in Algorithm A.6, which is the Algorithm 1 of Foster et al. (2021) combined with our GP prior. We summarize the comparison between DAD and our main safe AL training (Algorithm 2): (i) safety measurements are not present in DAD, (ii) the policy of DAD does not take the budget variable while $T_{\text{sim}}=T$ is fixed during the training, and (iii) functions $f_{0},...,f_{N_{f,q}}$ are sampled but the DAD objective simulates the observations $Y_{\text{init}},y_{1:T}$ on $f_{0}$ only while $f_{1},...,f_{N_{f,q}}$ are contrastive samples for noise contrastive estimations. Note that in this paper, $N_{f,q}$ denotes the number of functions |$|\{(f,q)\}|$ per set of GP hyperparameters. Here, we use $N_{f,q}$ to denote the number of contrastive functions.
@@
This objective was derived for parametric models, i.e. $p(\text{model parameters}|\text{data})$ in contrast to our $p(f|\text{data})$. Please see Foster et al. (2021) for details.
Notice that a further work, Ivanova et al. (2021), extended DAD to more general scenarios, for example, allowing intractable $p(y|f)$. However, their proposed learning objectives require another NN mapping, $\{(\mathcal{D},f)\}\rightarrow\mathbb{R}$ if described in our notation, to help estimate the involved intractable distributions. Such a mapping is not applicable for nonparametric $f$, and the objectives in Ivanova et al. (2021) were again not designed for AL of nonparametric functions. This is a different direction to what we need in this paper.
@@ C.3 Objectives: Safe AL
Figure A.C.6: Simulated safe AL objectives. $T_{\text{sim}}=1$. As described in Figure A.C.5, orange and green are $y,z$ and the corresponding GP predictions, while black curves are unconstrained objectives $\mathcal{H},\mathcal{H}_{\text{mean}}$. Our safe AL objectives (red) decorate the unconstrained objectives with our main min unsafe log probability (bottom, $\mathcal{S}_{\cdot}$q. 7) or with the appendix max safe log probability (top, $\mathcal{S}_{\cdot,\text{division}}$q. A.17). The objectives illustrated are meant to be maximized. Red dashed lines show the objectives if the safety probability is not clamped by $\gamma$. Note again that $y,z$ are available because our training scheme simulates AL before utilizing the evaluations for objective computations.
@@
In our main paper, we propose a safe AL objective
[…]
@@ D.1 Fourier Feature Functions
In Algorithm 3, the GP functions $f_{0}\sim\mathcal{GP}(0,k_{\theta}),q_{0}\sim\mathcal{GP}(0,k_{q})$ are approximated by Fourier features, which means each function sample is a linear combination of cosine functions (Rahimi & Recht, 2007; Wilson et al., 2020):
@@
$\omega_{i}\sim\mathcal{N}(0,\sqrt{v}^{2}),\bm{a}_{i}\sim\mathcal{N}(0,\text{diag}\{\bm{l}\}^{-1})$ are kernel dependent and $b_{i}\sim\text{Uniform}(0,2\pi)$. Each function has $L*(D+2)$ parameters. Larger $L$ lead to better approximations. An error bound w.r.t. $D$ and $L$ can be seen in Rahimi & Recht (2007). We set $L=100$.
@@
In our main paper, the analytical mean of window $\mathcal{X}=[0,1]^{D}$ is computed such that all functions can be shifted to zero mean in the particular domain. The analytical mean is the integral of $f_{0}(\bm{x}),q_{0}(\bm{x})$ divided by volume of $[0,1]^{D}$. We give examples for the one and two dimensional cases:
@@
The complexity of computing such mean is $\mathcal{O}\left(L2^{D}\right)$, and this can be distributed to time or space. With common GP and RBF kernel dimensions, e.g. $D\leq 10$, this complexity is negligible.
@@ D.2 Prior Mean of Safety Functions
When we sample $q=\mu_{q}+q_{0}-\mathbb{E}_{\bm{x}\in\mathcal{X}}[q_{0}(\bm{x})],q_{0}\sim\mathcal{GP}(0,k_{q})$ (see Algorithm 3), we aim for mean $\mathbb{E}[q]$ around 00 and variance of $q$ around $1$, as our deployment problems. We distribute the variance by making sure $c^{2}+v+\sigma_{q}^{2}$ is around $1$ (kernel variance $v$, observation noise level $\sigma_{q}$), in particular, we set $c^{2}=0.5$ while the kernel and the noise take the remaining half. The kernel and the noise are the main sources of functional stochasticity. The assumption of having additive variances is in fact not true for our $\mu_{q}$, but this setting is enough for the NN to learn. One may also consider the overall function variance as an amplitude of the function. The shape parameters $\bm{w}=(w_{1},...,w_{D})$ are sampled around value $20$, such that $\mu_{q}$ has around 10% to 100% of the space above safety threshold 00. We will summarize the sampling setting later in Appendix F.
@@ Appendix E Training Complexity
The NN forward passes takes $\mathcal{O}\left(\sum_{t=1}^{T}(N_{\text{init}}+t-1)^{2}\right)$ in time, as self attention has square complexity (Figure A.B.4Vaswani et al., 2017). The space complexity depends on the number of NN parameters.
@@ F.1 Policy Training
The batch sizes we set are listed in Table A.E.2. For each training pipeline, we train with a few different seeds. The optimizer is RAdam (Liu et al., 2020a). We set a lr scheduler to discount the lr by $2\%$ every $50$ training steps. We usually train with $200*50=10000$ steps, but we give the DAD baseline more steps to converge (still not competitive on GP functions when $D\geq 2$). The exact traininig steps and times are summarized in Table A.F.5. Each training job was conducted on one NVIDIA RTX3090 (24GB GPU memory).
@@
For the DAD baseline (described in Section C.2Algorithm A.6), we take the code from Foster et al. (2021); Ivanova et al. (2021). The DAD implementation utilizes Pyro (Bingham et al., 2018) to condition the same sequence of queries on different functions, which however involve CPUs for each loss computation. In comparison, our loss functions are computed completely with PyTorch which requires little CPU-GPU interaction. This is why the training time per step of DAD is not necessarily faster than our objectives (Table A.F.5), despite DAD time complexity is much smaller (Table A.E.3).

### v2

@@ C.1.2 Objectives: unconstrained AL - regularized entropy
Here, we discuss empirical issues of unconstrained entropy and our approach using mutual information objectives [17, 23].
Maximizing an entropy objective favors a set of distinctive points, which naturally encourages points at the border, as they are the most scattered. Guestrin et al. [17] argued that exploring the border may be suboptimal because we would rather explore inside the space. The authors further proposed a mutual information acquisition criterion to tackle this problem, at least in conventional AL settings. In our training scheme, the objectives $\mathcal{H}(\phi)$ and $\mathcal{H}_{\text{mean}}(\phi)$ sometimes overemphasize the border and completely ignore the inner region of $\mathcal{X}$, which is clearly not desired. Part of the reasons can be the tanh layer in our NN (Figure A.B.4), which, although helping us bound the NN outputs in $\mathcal{X}$, has vanishing gradients at the boundary. To tackle this issue, we wish to derive a mutual information objective suitable for our framework.
In active GP learning, Guestrin et al. [17] define a mutual information criterion as the reduction of entropy in an unexplored region: $\mathbb{H}(y(\bm{x}_{\phi,1}),...,y(\bm{x}_{\phi,T})|Y_{\text{init}})-\mathbb{H}(y(\bm{x}_{\phi,1}),...,y(\bm{x}_{\phi,T})|Y_{\text{init}},\hat{y}(\mathcal{X}\setminus\bm{X}_{\phi}))$, $\hat{y}(\mathcal{X}\setminus\bm{X}_{\phi})$ means the output estimates corresponding to $\mathcal{X}\setminus\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\}$. The original paper considered small discrete $\mathcal{X}$ where $\hat{y}(\mathcal{X}\setminus\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\})$ is a computable set of variables. In our framework, $\mathcal{X}$ is a continuous space, and thus this is not well-defined. Even if $\mathcal{X}$ is discrete, conditioning on a large pool (fine discretization) is computationally heavy, i.e. GP cubic complexity $\mathcal{O}\left(|\mathcal{X}|^{3}\right)$ (Appendix A). A discrete pool also enforces a classifier-like policy $\phi$, as the policy needs to select points from a pool, which prohibits us from utilizing the current NN structure developed on top of Foster et al. [11], Ivanova et al. [20].
To derive mutual information objectives suitable for our learning framework, note that entropy reduction is a regularized entropy objective. We propose a simple yet effective approach: compute the regularization term only on a sparse set of $N_{\text{grid}}$ samples $\left(\bm{X}_{\text{grid}},Y_{\text{grid}}\right)\subseteq\mathcal{X}\times\mathcal{Y}$. Then we turn the acquisition functions proposed by Guestrin et al. [17], Krause & Guestrin [22] into the following training objectives
@@
$\begin{split}\mathcal{I}_{\text{mean}}(\phi)=&\mathbb{E}\left[\frac{\mathbb{H}(y(\bm{x}_{\phi,1:T_{\text{sim}}})|Y_{\text{init}})-\mathbb{H}(y(\bm{x}_{\phi,1:T_{\text{sim}}})|Y_{\text{init}},Y_{\text{grid}})}{N_{\text{init}}+T_{\text{sim}}}\right].\end{split}$ (A.16)
The expectation is over GPs ($\theta,\sigma^{2}$) and AL functions ($f,T_{\text{sim}},\mathcal{E}_{\text{init}},\epsilon_{1:T_{\text{sim}}}$). $N_{\text{grid}}$ should be much larger than $T$. Maximizing these objectives encourages $\{\bm{x}_{\phi,1},...,\bm{x}_{\phi,T}\}$ to track subsets of $\bm{X}_{\text{grid}}$. The intuition is two-fold: (i) we can view them as entropy objectives regularized by an additional search space indicator, or (ii) we can view them as imitation objectives because a subset of grid points, if happens to have small joint likelihood or large joint entropy, maximizes the objective.
@@ C.2 Objectives: DAD baseline
1:2.1, $T$, $N_{\text{init}}$
2:draw a batch of $(\theta,\sigma^{2})$
3:for each $(\theta,\sigma^{2})$ do
4:     draw $f_{0,\text{raw}}\sim\mathcal{GP}(0,k_{\theta})$
5:     $f_{0}(\cdot)=f_{0,\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{0,\text{raw}}(\bm{x})]$
6:     draw $\bm{X}_{\text{init}}\sim\text{Uniform}[\mathcal{X}],|\bm{X}_{\text{init}}|=N_{\text{init}}$
7:     draw $\mathcal{E}_{\text{init}}\sim\mathcal{N}(0,\sigma^{2}I),Y_{\text{init}}=f_{0}(\bm{X}_{\text{init}})+\mathcal{E}_{\text{init}}$
8:     $\mathcal{D}_{0}=\{\bm{X}_{\text{init}},Y_{\text{init}}\}$
9:     for $t=1,...,T$ do
10:         $\bm{x}_{\phi,t}=\phi(\mathcal{D}_{t-1})$
11:         draw $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}),y_{\phi,t}=f_{0}(\bm{x}_{\phi,t})+\epsilon_{t}$
12:         $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{\phi,t},y_{\phi,t}\}$
13:     end for
14:     draw $f_{l,\text{raw}}\sim\mathcal{GP}(0,k_{\theta}),l=1,...,N_{f,q}$
15:     $f_{l}(\cdot)=f_{l,\text{raw}}(\cdot)-\mathbb{E}_{\bm{x}\in\mathcal{X}}[f_{l,\text{raw}}(\bm{x})],l=1,...,N_{f,q}$
16:end for
17:compute DAD loss (Eq. A.17), update $\phi$
We describe the DAD baseline in our GP context [11]. DAD is a scheme which samples tasks and learns to perform Bayesian optimal experimental design originally aiming at parametric models. Bayesian optimal experimental design aims to query data to learn about a posterior Bayesian model [36]. In a GP learning problem, this means collecting $\bm{x}_{1:T},y_{1:T}$ to make inference $p(f(\cdot)|\bm{X}_{\text{init}},Y_{\text{init}},\bm{x}_{1:T},y_{1:T})$. DAD follows similar training procedure: we (i) sample tasks from a prior, and (ii) learn a data querying policy by optimizing the DAD objective. To train with DAD, one may use our main training algorithm because we generate more quantities than needed by the DAD objective. For clarity, however, we write down a clean version of DAD training in Algorithm A.6, which is the Algorithm 1 of Foster et al. [11] combined with our GP prior. We summarize the comparison between DAD and our main safe AL training (Algorithm 2): (i) safety measurements are not present in DAD, (ii) the policy of DAD does not take the budget variable while $T_{\text{sim}}=T$ is fixed during the training, and (iii) functions $f_{0},...,f_{N_{f,q}}$ are sampled but the DAD objective simulates the observations $Y_{\text{init}},y_{1:T}$ on $f_{0}$ only while $f_{1},...,f_{N_{f,q}}$ are contrastive samples for noise contrastive estimations. Note that in this paper, $N_{f,q}$ denotes the number of functions |$|\{(f,q)\}|$ per set of GP hyperparameters. Here, we use $N_{f,q}$ to denote the number of contrastive functions.
@@
This objective was derived for parametric models, i.e. $p(\text{model parameters}|\text{data})$ in contrast to our $p(f|\text{data})$. Please see Foster et al. [11] for details.
Notice that a further work, Ivanova et al. [20], extended DAD to more general scenarios, for example, allowing intractable $p(y|f)$. However, their proposed learning objectives require another NN mapping, $\{(\mathcal{D},f)\}\rightarrow\mathbb{R}$ if described in our notation, to help estimate the involved intractable distributions. Such a mapping is not applicable for nonparametric $f$, and the objectives in Ivanova et al. [20] were again not designed for AL of nonparametric functions. This is a different direction to what we need in this paper.
@@ C.3 Objectives: safe AL
Figure A.C.6: Simulated safe AL objectives. $T_{\text{sim}}=1$. As described in Figure A.C.5, orange and green are $y,z$ and the corresponding GP predictions, while black curves are unconstrained objectives $\mathcal{H},\mathcal{H}_{\text{mean}}$. Our safe AL objectives (red) decorate the unconstrained objectives with our main min unsafe log probability (bottom, $\mathcal{S}_{\cdot}$q. 8) or with the appendix max safe log probability (top, $\mathcal{S}_{\cdot,\text{division}}$q. A.18). The objectives illustrated are meant to be maximized. Red dashed lines show the objectives if the safety probability is not clamped by $\gamma$. Note again that $y,z$ are available because our training scheme simulates AL before utilizing the evaluations for objective computations.
@@
In our main paper, we propose a safe AL objective (Eq. 8)
$\mathcal{S_{H}}(\phi)$ $=\mathbb{E}\left[\frac{-\log p(y_{\phi,1},...,y_{\phi,T_{\text{sim}}}|Y_{\text{init}})-\sum_{t=0}^{T_{\text{sim}}-1}\log\ \text{max}(\gamma,p(z(\bm{x}_{\phi,t+1})<0|z_{\phi,1:t},Z_{\text{init}}))}{N_{\text{init}}+T_{\text{sim}}}\right].$
@@ D.1 Fourier feature functions
In Algorithm 3, the GP functions $f_{\text{raw}}\sim\mathcal{GP}(0,k_{\theta}),q_{\text{raw}}\sim\mathcal{GP}(0,k_{q})$ are approximated by Fourier features, which means each function sample is a linear combination of cosine functions [35, 53]:
@@
$\omega_{i}\sim\mathcal{N}(0,\sqrt{v}^{2}),\bm{a}_{i}\sim\mathcal{N}(0,\text{diag}\{\bm{l}\}^{-1})$ are kernel dependent and $b_{i}\sim\text{Uniform}(0,2\pi)$. Each function has $L*(D+2)$ parameters. Larger $L$ lead to better approximations. An error bound w.r.t. $D$ and $L$ can be seen in Rahimi & Recht [35]. We set $L=100$.
@@
In our main paper, the analytical mean of window $\mathcal{X}=[0,1]^{D}$ is computed such that all functions can be shifted to zero mean in the particular domain. The analytical mean is the integral of $f_{\text{raw}}(\bm{x}),q_{\text{raw}}(\bm{x})$ divided by volume of $[0,1]^{D}$. We give examples for the one and two dimensional cases:
@@
The complexity of computing such mean is $\mathcal{O}\left(L2^{D}\right)$, and this can be distributed to time or space. With common GP and RBF kernel dimensions, e.g. $D\leq 5$ or $10$, this complexity is negligible.
@@ D.2 Prior mean of safety functions
When we sample $q=\mu_{q}+q_{\text{raw}}-\mathbb{E}_{\bm{x}\in\mathcal{X}}[q_{\text{raw}}(\bm{x})],q_{\text{raw}}\sim\mathcal{GP}(0,k_{q})$ (see Algorithm 3), we aim for mean $\mathbb{E}[q]$ around 00 and variance of $q$ around $1$, as our deployment problems. We distribute the variance by making sure $c^{2}+v+\sigma_{q}^{2}$ is around $1$ (kernel variance $v$, observation noise level $\sigma_{q}$), in particular, we set $c^{2}=0.5$ while the kernel and the noise take the remaining half. The kernel and the noise are the main sources of functional stochasticity. The assumption of having additive variances is in fact not true for our $\mu_{q}$, but this setting is enough for the NN to learn. One may also consider the overall function variance as an amplitude of the function. The shape parameters $\bm{w}=(w_{1},...,w_{D})$ are sampled around value $20$, such that $\mu_{q}$ has around 10% to 100% of the space above safety threshold 00. We will summarize the sampling setting later in Appendix F.
@@ Appendix E Training complexity
The NN forward passes takes $\mathcal{O}\left(\sum_{t=1}^{T}(N_{\text{init}}+t-1)^{2}\right)$ in time, as self attention has square complexity (Figure A.B.4[52]). The space complexity depends on the number of NN parameters.
@@ F.1 Policy training
The batch sizes we set are listed in Table A.E.2. For each training pipeline, we train with a few different seeds. The optimizer is RAdam [28]. We set a lr scheduler to discount the lr by $2\%$ every $50$ training steps. We usually train with $200*50=10000$ steps, but we give the DAD baseline more steps to converge (still not competitive on GP functions when $D\geq 2$). The exact traininig steps and times are summarized in Table A.F.5. Each training job was conducted on one NVIDIA RTX3090 (24GB GPU memory).
NN architectural hyperparameters:
The NN configurations are taken largely from Ivanova et al. [20]: The transformer block is taken as it is; the number of hidden neurons in the MLPs (Data Embed, Budget Feed Forward, and Decision Feed Forward) are all set to 512. The remaining parameter is the embedding dimension (the dimension of $e_{t},E_{t},ES_{t},EB$, and the input dimension of the Decision Feed Forward), which determines the size of the transformer layers and plays a crucial role in the NN size. In our main paper, we set this dimension to 128, but we provide an ablation study on this dimension in Appendix G.
@@
For the DAD baseline (described in Section C.2Algorithm A.6), we take the code from Foster et al. [11], Ivanova et al. [20]. The DAD implementation utilizes Pyro [5] to condition the same sequence of queries on different functions, which however involve CPUs for each loss computation. In comparison, our loss functions are computed completely with PyTorch which requires little CPU-GPU interaction. This is why the training time per step of DAD is not necessarily faster than our objectives (Table A.F.5), despite DAD time complexity much smaller (Table A.E.3).